OpenStudy (anonymous):

captain snorkel is supervising the construction of a new sub. the design calls for the installation of 20 round view ports with each view port being 19 inches in diameter. how many square feet of transparent materal will be included in the view ports?

OpenStudy (anonymous):

19 inches= 19/12 ft radius= diameter/2= 19/24 ft Area 1 view port= pi* radius^2= pi* (19/24)^2 20 view ports, total area is= 20*pi*(19/24)^2= 29.379 ft^2 <--answer

OpenStudy (anonymous):

39.379 actually